# Learning Machine Morality through Experience and Interaction

### v2

@@ 2.2 Bottom-up Approaches
Given the disadvantages of top-down approaches to the design of moral agents outlined above, an alternative methodology may be needed for developing machine morality. Many recent advances in other areas of Artificial Intelligence can be attributed to moving away from hand-engineered rule-based algorithms and features towards more flexible learning systems based on deep learning (?) - for example in vision (?) and natural language processing (?, ?). An increasingly powerful framework for developing systems for decision-making in particular is RL, where an agent learns by trial-and-error while interacting with an environment and receiving a reward signal (?) (see Section 3.1 for a detailed definition). Of particular relevance to social capabilities in AI, RL has allowed researchers to develop super-human game playing agents (?, ?, ?). More generally, ? (?) argue that the pursuit of reward may be sufficient for all kinds of intelligence to arise. Given the success of learning over rule-based algorithms across these domains, a promising direction for ethical reasoning may be to enable agents to develop a general moral intuition bottom-up (?), by learning from interactions with a social environment over time (?). A seminal paper proposing the use of RL for ethical learning is ? (?). Focusing on single-agent ethical dilemmas, the authors outlined the strengths of RL over rule-based or Bayesian approaches. This RL-based approach would also align with the view of some developmental psychologists, who believe that humans learn morality gradually, starting with simple norms in early childhood and advancing to complex ethical principles later on (?, ?, ?).
@@
Specific examples of bottom-up moral learning can include Supervised Learning on a data set of scenarios (e.g., dilemmas), in which a set of actions in response to a given situation are labeled as morally permissible or not, as in ? (?). More recently, with the emergence of Large Language Models (LLMs), another bottom-up approach to representing morality in systems could be to train LLMs on morally relevant text - such as moral philosophical writings, fables or religious texts. Within the RL domain, one example of fully bottom-up learning is Inverse RL, where a reward function is learned directly from expert (e.g., human) demonstrations of desired behavior (?). Inverse RL for morality in particular has been explored by ? (?, ?, ?). Extensions of this include (Inverse) Multi-Objective RL (?) in situations in which multiple humans might contradict one another in terms of moral (and other) norms (for example, see work by ?). A more recent yet related approach is Reinforcement Learning from Human Feedback (RLHF), in which humans rank outputs from a pre-trained model to display their preferences, and these rankings are used as a signal to fine-tune the model (?, ?). The strength of RLHF that has contributed to its recent growth in popularity is that it allows researchers to use relative rankings to infer implicit values in concepts that may otherwise be difficult to formalize as rewards (indeed, as argued in Section 2.1, many moral preferences may not be possible to formalize in an explicit way). This has been demonstrated by the successful reduction in the toxicity of language model outputs following fine-tuning with RLHF, without having to explicitly define what toxicity means in all possible linguistic constructions (?). Alternatively, if one has control over the environment as well as the agent, social mechanisms can be implemented that may lead agents to learn moral policies on their own - examples of this have included implementations of mechanisms inspired by Rawls’s (?) idea of the Veil of Ignorance (see ?), or a partner selection mechanism in social dilemmas (?).
@@
A more general advantage of a learning agent over a rule-based one is that agents that learn continuously are able to adapt to the potentially changing dynamics or evolving morality of a given society. Additionally, learned policies which utilize function approximation, for example via Deep Reinforcement Learning (?), have the potential to generalize to unseen situations, or ‘states’ in RL, and will try to learn a policy that accommodates potentially different or even contradicting sets of preferences without causing deadlock - for an example, see ? (?).
@@
In addition, Railton in (?) argues that AI may be more robust to potential adversarial attacks if ethical reasoning forms a part of an agent’s monolithic training from the start, with a root connection to the agent’s other learned policies, rather than a top-down add-on which can easily be removed without reduction in other intelligent capacities of the AI system. Indeed, such a developmental connection between ethical reasoning and other abilities such as problem-solving exists in humans, as observed in lesion studies (?). Interestingly, it has recently been suggested that fine-tuned LLMs may fall victim to this exact problem at the end of training - ? (?) suggest that fine-tuning model weights may merely create a ‘safety wrapper’ on otherwise unaltered core models, making the effects of safety-focused fine-tuning easily reversible. This suggests that methods of learning morality that alter the model itself from the start may be preferred for safety-critical applications.
@@
More generally, despite their strengths, full bottom-up learning methods including RL may also present a set of disadvantages for the development of machine morality. First of all, RL tends to be sample inefficient - even in simple environments, agents require a large amount of experience to learn the optimal policy. This is usually addressed by initially training agents in simulation, in an artificial environment. However, this might require the development of extremely complex simulated social environments, which might be very difficult to define in practice. The problem of coverage is key in this context.
@@
Furthermore, learning agents are not guaranteed to learn exactly the values or behaviors intended by their designer - a problem known broadly as misalignment. For example, safety researchers ? (?) and ? (?) suggest that RL-based agents may potentially perform the so-called ‘reward-hacking’, in which they display behavior that leads to the optimal cumulative reward received on the test set of problems, even if the underlying value function is not actually encoding truly safe policies. A related issue is problem misspecification, which occurs when the simulated environment is not designed to correctly reflect the intended true setting, so policies learned in simulation do not result in desirable behavior in the real world (?, ?, ?). Notably, neither reward-hacking, nor problem misspecification are solved by the increasingly popular RLHF methodology, as discussed by ? (?).
@@
A further issue that can arise for learning agents is data or environment poisoning - a type of adversarial attack that arises when a designer of an AI system has no full control of the data sets used in pre-training. For example, in classification tasks in the computer vision domain (?), if an adversary implants a small area of ‘trigger’ pixels into a few data points, it can cause an otherwise well-performing model to learn false associations and thus misclassify samples containing this trigger at inference time. Poisoning attacks like these are a risk for any general learning agents, but in relation to moral or social agents in particular, such attacks may lead to explicit harms towards other agents or humans in the environment, thus posing a more significant safety risk.
@@
Thus, the behavior of RL agents must be carefully evaluated in different kinds of scenarios to ensure the learned representations are aligned with the designer’s intended moral values as observed in a variety of situations. We discuss this further in Section 4. An added issue in RLHF in particular is misalignment among the humans who provide the ranking data (?) - so the choice of human sample to collect feedback from becomes a crucial one, and cultural differences are likely to have a strong influence on the moral choices made by the humans and thus learned by the fine-tuned model (?).
@@
Finally, traditional RL faces specific issues when applied in particular to multi-agent decision-making situations, such as social dilemmas (? - see Section 3.4.2 for a more detailed discussion), in which every agent must make a trade-off between individual and social benefit from their actions in the long term. Here, we observe that due to the social or moral uncertainty (?) and the presence of numerous learning agents in an environment, traditional ‘selfish’ RL agents trained to maximize their own game reward often learn inefficient (i.e., defective) policies or norms (?, ?, ?). Humans playing the same dilemma games manage to find ways to cooperate, as demonstrated by Behavioral Game Theory (?).
@@ 2.3 Hybrid Approaches: Pragmatic Solutions?
Given the considerations above, Wallach and colleagues (?, ?) have recommended a hybrid approach for developing AI morality, which combines some sort of formal top-down definition of moral principle with bottom-up learning. We support this view, especially in presence of learning and adaptive AI systems. As discussed, fully bottom-up approaches do not often provide guarantees and they are generally not preferable for potentially risky situations, such as in human-robot interaction or financial applications. Top-down approaches, on the other hand, are not appropriate when the principles and constraints are difficult to specify or they are dynamic.
Despite the promise of a hybrid approach in many scenarios, very little existing work has used it for designing ethical learning agents. In this paper, we argue that learning through experience and interaction with explicit moral principles provides an pragmatic way to implementing AI agents that learn morality in a way that is safe and flexible. For example, variations of RL can embed a variety of moral principles top-down as objective or reward functions, while simultaneously allowing agents to learn from experience how to best behave according to these principles in a potentially changing or uncertain social environment (for a parallel discussion in RL applications with human advice outside the domain of ethics, see ?). In particular, RL approaches which use top-down principles to modify the reward structure or the environment can embed the understanding of morality which has been developed over centuries in different domains such as Philosophy, Psychology, Biology and Law, to name a few.
@@
In the subsequent section, we formalize the learning problem for RL agents, and consider three examples of how moral principles can be represented. The most ‘controlled’ or top-down of these is RL with safety constraints (for example, using ‘shields’ to avoid harm in robotics tasks or self-driving cars - ?; or constraining RL by a deontic-logic normative supervisor - ?). This case is not fully top-down since learning is present, but the moral aspect of the problem is hard-coded, so we place them towards the top-down end of the continuum in Figure 1. The least constrained hybrid example we present here is Constitutional AI (?) - a technique for fine-tuning LLMs with RL using feedback from a ‘constitution’ of AI models, each of which is prompted to follow certain explicit principles. Finally, between the control of Safe RL and the complexity of LLM fine-tuning is a methodology implementing RL with intrinsic moral rewards (?, ?, ?). Using social dilemma scenarios (?, ?) to illustrate the intrinsic rewards approach, we also overview possible moral frameworks that could be encoded as reward functions in these games in general, and provide a few concrete examples. Throughout the case studies, we refer to existing experimental results that show that these RL-based hybrid approaches can successfully lead to the design of agents that learn moral policies in social environments.
3 Designing Moral Agents with Reinforcement Learning
@@ 3.1 Background
Reinforcement Learning (RL) is a method that directly models decision-making, as opposed to prediction or classification - as in Supervised or Unsupervised Learning. In the RL framework, learning occurs incrementally over discrete time steps $t\in T$. At every time step $t$, an agent observes a state of the environment $s^{t}$ and takes an action $a^{t}$ using a policy (a mapping between states and actions). The environment then returns a reward $R^{t+1}$ and a next state $s^{t+1}$. Over a number of iterations of such interactions with an environment, an agent learns to take actions that maximize the discounted future reward (discounted to make rewards received further in the future less important for the present calculations). Frequently, RL implementations use the Markov Decision Process framework, in which all necessary information for the agent to make a decision is contained within the last available state.
@@
Agents use an RL algorithm to learn what actions are best to take to maximize their cumulative reward. In this paper, we focus on the set of approaches known as model-free, where an agent does not know the model of how the environment may respond to their actions - we believe that this approach applies intuitively to uncertain social situations where other agents whose behavior is not directly predictable constitute a part of the environment. One popular model-free RL algorithm is Q-Learning, introduced by ? (?), where an agent learns the ‘value’ of each possible state-action pair $Q(s,a)$ (i.e., the estimated discounted cumulative reward an agent may expect to receive by taking this action in this state, and assuming they act optimally from then on).
@@
Two versions of Q-learning exist: a tabular version for small state spaces and Deep Q-Network, or DQN (?) for value function approximation in state spaces too large to learn directly (see Appendix A for a detailed formulation). As shown in Figure 2, in tabular Q-learning, the learned value function constitutes a table mapping every state-action pair to the corresponding Q-value. In DQN, on the other hand, the learned representation is a tuned neural network which approximates the state-action value function (?), which takes a state vector representation as input and returns the relative value of taking each possible action as an output. It is this representation that allows DQN agents to generalize knowledge across states, enabling them to choose actions based on past experience even if they have not seen this particular state before.
@@
Figure 2: The representations learned in Q-learning RL algorithms. In tabular Q-learning - left (?), an agent learns a direct mapping - in the form of a table - from every state to the value of taking each possible action. In DQN - right (?), an agent trains a neural network which takes a representation of the state as input and outputs the values of taking each possible action. DQN uses function approximation, which means that not all states need to be observed, because the trained neural network can generalize across states and is able to output a value for a state not seen during prior training.
@@
These representations and the learned value functions can then be used to choose actions according to a certain policy $\pi$ for example, using an $\epsilon$-greedy policy (Equation 1) an agent would take a ‘greedy’ action most of the time, i.e., the action with the greatest estimated Q-value available from this state, but act randomly $\epsilon\%$ of the time, sampling uniformly at random from the available action space $A$. Our formulation uses an online RL paradigm where the agent is evaluated as it continuously learns to adapt to its social environment. The random exploration in this policy allows for new states to be explored by the agent, an essential step in learning via online RL.
@@
Most morally relevant environments are likely to be multi-agent. In these cases, multiple players learn (i.e., iteratively update their Q-value estimates) in parallel, and, for each agent, their opponent(s) constitute a part of the environment. The state for a player is then defined as a history of their opponents’ previous moves. We will now explore three examples of the use of RL for learning morality via explicitly defined principles.

### v3

@@ 2.2 Bottom-up Approaches
Given the disadvantages of top-down approaches to the design of moral agents outlined above, an alternative methodology may be needed for developing machine morality. Many recent advances in other areas of Artificial Intelligence can be attributed to moving away from hand-engineered rule-based algorithms and features, towards more flexible and adaptive systems based on deep learning (?). Particularly prominent examples of the power of bottom-up learning come from the domains of vision (?) and natural language processing (?, ?). In the domain of developing systems for decision-making in particular, an increasingly powerful framework for is RL, where an agent learns by trial-and-error while interacting with an environment and receiving a reward signal (?; see Section 3.1 below for a detailed definition). Of particular relevance to social capabilities in AI, RL has allowed researchers to develop game playing agents able to perform at super-human level (?, ?, ?). More generally, ? (?) argue that the pursuit of reward may be sufficient for all kinds of intelligence to arise in artificial systems. Given the success of learning over rule-based algorithms across these domains, a promising direction for ethical alignment may be to enable agents to develop a general moral intuition from the bottom-up (?), by learning from interactions with a social environment over time (?). A seminal paper proposing the use of RL for ethical learning is ? (?), in which, focusing on single-agent ethical dilemmas, the authors outline the strengths of RL over rule-based or Bayesian approaches. This RL-based approach might also align better than rule-based approaches with the view of some developmental psychologists, who believe that humans learn morality gradually, starting with simple norms in early childhood and advancing to complex ethical principles later on (?, ?, ?, ?).
@@
Specific examples of bottom-up moral learning can include Supervised Learning on a data set of scenarios (e.g., dilemmas), in which a set of actions in response to a given situation are labelled as morally permissible or not, as in ? (?). More recently, with the emergence of LLMs, another bottom-up approach to representing morality in systems could be to train LLMs on morally relevant text - such as moral philosophical writings, fables or religious texts. Within the RL domain, one example of fully bottom-up learning is Inverse RL, where a reward function is learned directly from expert (e.g., human) demonstrations of desired behaviour (?, ?). Inverse RL for morality in particular has been explored by ? (?, ?) and ? (?). Extensions of this, which aim to address the pluralistic nature of human moral preferences, include (Inverse) Multi-Objective RL (?) in situations in which multiple humans might contradict one another in terms of moral (and other) norms - for example, see work by ? (?). A more recent yet related approach is Reinforcement Learning from Human Feedback (RLHF), in which samples of humans rank outputs from a pre-trained model to display their preferences, and these rankings are used as a signal to fine-tune the model (?, ?). The strength of RLHF that has contributed to its recent growth in popularity is that it allows researchers to use relative rankings to infer implicit values in concepts that may otherwise be difficult to formalise as rewards (indeed, as argued in Section 2.1, many moral preferences may not be possible to formalise in an explicit way). This has been demonstrated by the successful reduction in the toxicity of language model outputs following fine-tuning with RLHF, without having to explicitly define what toxicity means in all possible linguistic constructions (?). Alternatively, if one has control over the environment as well as the agent, social mechanisms can be implemented that may lead RL agents to learn moral policies on their own - examples of this have included implementations of mechanisms inspired by Rawls’s idea (?) of the Veil of Ignorance (see ?), or a partner selection mechanism in social dilemmas (?).
@@
A more general advantage of a learning agent over a rule-based one is that agents that learn continuously are able to adapt to the potentially changing dynamics or evolving morality of a given society. Additionally, learned policies which utilise function approximation, for example via Deep Reinforcement Learning (?), have the potential to generalise to unseen situations, or ‘states’ in RL. Furthermore, a deep RL-based system may be able to learn a policy that accommodates potentially different or even contradicting sets of preferences without causing deadlock (e.g., ?).
@@
In addition, ? (?) argues that AI may be more robust to potential adversarial attacks if ethical reasoning forms a part of an agent’s monolithic training from the start, with a root connection to the agent’s other learned policies, rather than a top-down add-on which can easily be removed without reduction in other intelligent capacities of the AI system. Indeed, such a developmental connection between ethical reasoning and other abilities such as problem-solving exists in humans, as observed in lesion studies (?). Interestingly, it has recently been suggested that fine-tuned LLMs may fall victim to this exact problem at the end of training - ? (?) suggest that fine-tuning model weights may merely create a ‘safety wrapper’ on otherwise unaltered core models, making the effects of safety-focused fine-tuning easily reversible. This suggests that methods of learning morality that influence the model earlier on (i.e., in pre-training) may be preferred for safety-critical applications.
@@
More generally, despite their strengths, full bottom-up learning methods including RL may also present a set of disadvantages for the development of machine morality. First of all, RL tends to be sample inefficient - even in simple environments, agents require a large amount of experience to learn the optimal policy. This is usually addressed by initially training agents in simulation, in an artificial environment. However, for a domain as abstract as moral alignment this might require the development of extremely complex simulated social environments, which might be very difficult to define in practice. The problem of coverage is key in this context. A related issue is problem misspecification, which occurs when the simulated environment is not designed to correctly reflect the intended true setting, so policies learned in simulation do not result in desirable behaviour in the real world at deployment time (?, ?, ?).
@@
Learning agents are also not guaranteed to learn exactly the values or behaviours intended by their designer - a problem known broadly as misalignment. Safety researchers ? (?) and ? (?) suggest that RL-based agents may potentially perform the so-called ‘reward-hacking’, in which they display behaviour that obtains the optimal cumulative reward on the test set of problems, even if the underlying value function is not actually encoding truly safe policies. Notably, neither reward-hacking, nor problem misspecification are solved by the increasingly popular RLHF methodology, as discussed by ? (?).
@@
A further issue that can arise for learning agents is data or environment poisoning - a type of adversarial attack that arises when a designer of an AI system does not have full control of the data sets used in pre-training. For example, in classification tasks in the computer vision domain (?), if an adversary implants a small area of ‘trigger’ pixels into a few training data points, it can cause an otherwise well-performing model to learn false associations and thus misclassify samples containing this often unnoticeable trigger at inference time. Poisoning attacks like these are a risk for any general learning agents, but in relation to moral or social agents in particular, such attacks may lead to explicit harms towards other agents or humans in the environment, thus posing a more significant safety risk.
@@
Thus, the behaviour of RL agents must be carefully evaluated in different kinds of scenarios to ensure the learned representations are aligned with the designer’s intended moral values as observed in a variety of situations. We discuss the problem of agent evaluation further in Section 4. An added issue in human feedback-based RL in particular (i.e., RLHF) is misalignment among the humans who provide the ranking data (?) - therefore, the choice of human sample to collect feedback from becomes a crucial one, and cultural differences are likely to have a strong influence on the moral choices made by the humans and thus learned by the fine-tuned model (?, ?).
@@
Finally, traditional RL faces specific issues when applied in particular to multi-agent decision-making situations (?), such as social dilemmas, in which every agent must make a trade-off between individual and social benefit from their actions in the long term (? - see Section 3.4.2 for a more detailed discussion). Here, we observe that due to the social or moral uncertainty (?) and the presence of numerous learning agents in an environment, traditional ‘selfish’ RL agents trained to maximise their own game reward often learn inefficient (i.e., defective) policies or norms (?, ?, ?). Humans playing the same dilemma games manage to find ways to cooperate, as demonstrated by Behavioural Game Theory (?).
@@ 2.3 Hybrid Approaches: Pragmatic Solutions?
Given the considerations above, Wallach and colleagues (?, ?) have recommended a hybrid approach for developing AI morality, which combines some sort of formal top-down definition of moral principles with bottom-up inference. We support this view, especially in terms of developing learning and adaptive AI systems.
As discussed, fully bottom-up approaches rarely provide guarantees and they are generally not preferable for potentially risky situations, such as in human-robot interaction or financial applications. A hybrid RL-based implementation can reduce risks such as reward-hacking by defining custom reward functions that more explicitly and directly encode the moral values of interest. For example, RL-based implementations can embed a variety of moral principles top-down as objective or reward functions (relying on the understanding of morality which has been developed over centuries in different domains such as Philosophy, Psychology, Biology and Law, to name a few), while simultaneously allowing agents to learn from experience how to best behave according to these principles in a potentially changing or uncertain social environment. Practical examples of this include RL with intrinsic rewards, which we review in Sections 3.4 and 3.5, and safety-constrained RL, described in Section 3.2. The sample inefficiency inherent in purely bottom-up RL can also be addressed by hybrid methods, whereby interim rewards or curricula can be designed to be more informative for an agent than potentially sparse environment rewards. Finally, explicitly identifying principles for training agents might help create training data that reduces the likelihood of misspecification, poisoning attacks or convergence to inefficient norms. A practical example of this is Constitutional AI, which we review as a case study in Section 3.3.
Top-down approaches, on the other hand, are not appropriate when the principles and constraints are difficult to specify, implicit or dynamic. Hybrid approaches, by bringing in the element of learning from experience, can allow agents to learn more general policies while still adhering to specific constraints (as in the example of RL with safety constraints, reviewed in Section 3.2), or implicit and multi-dimensional preferences present in the training data (as in the case of Constitutional AI, reviewed in Section 3.3). Additionally, learning signals, such as rewards, can be defined for agents in terms of outcomes rather than rules or constraints on the action space, allowing for more goal-directed flexible policies to be learned. An example of this is Utilitarian rewards in social dilemma games, reviewed in Section 3.4.3. Finally, continual learning approaches can allow agents to adapt to changes in the environment, such as changes in preferred behaviours and associated rewards (?). One example of this is the ability to ‘unlearn’ a previously developed strategy under new circumstances, as discussed in the work of ? (?), who first fine-tune LLM agents for a selfish strategy, and then for a prosocial moral one - we review this work in Section 3.5.
Despite the promise of a hybrid approach in many scenarios, very little existing work has used it for designing ethical learning agents. The case studies we present here map almost the entirety of the existing approaches in this space (?, ?). In this paper, we argue that learning through experience and interaction with explicit moral principles provides a pragmatic way to implementing AI agents that learn morality in a way that is safe and flexible.
@@
In the subsequent section, we formalise the moral learning problem for RL agents, and consider a set of specific examples of how moral principles can be represented and learned by a hybrid system. The most ‘controlled’ or top-down of these is RL with safety constraints, which has been applied to developing ‘shields’ to avoid harm in robotics tasks or self-driving cars (?), or constraining RL by a deontic-logic normative supervisor (?). The latter example is considered as a case study in Section 3.2. This case is not fully top-down since learning is present, but the moral aspect of the problem is hard-coded, so we place it towards the top-down end of the continuum in Figure 2. The least constrained hybrid example we present here is Constitutional AI (?) - a technique for fine-tuning LLMs with RL using feedback from a ‘constitution’ of AI models, each of which is prompted to follow certain explicit principles. This case study is considered in Section 3.3. Finally, between the control of Safe RL and the complexity of LLM feedback-based fine-tuning is a methodology that implements RL with intrinsic moral rewards (?, ?, ?). In Sections 3.4 and 3.5, we present two case studies where intrinsic rewards have been defined in social dilemma scenarios (?, ?) and applied to training pure-RL agents (?) as well as LLM-based ones (?). After presenting these two specific existing implementations, we also overview a larger set of possible moral frameworks that could be encoded as reward functions in these games in general. In aggregate, the case studies show that these RL-based hybrid approaches can successfully lead to the design of agents that learn moral policies in social environments.
3 Designing AI Agents based on Moral Principles
Moral AI agents can be designed with explicit moral principles using a variety of techniques. The examples presented in this paper centre around two domains: RL and LLM fine-tuning. In practice, moral fine-tuning of LLMs also often relies on RL. Examples include fine-tuning with intrinsic rewards (described in Section 3.5) or with reward based on feedback from a constitution of other LLMs (as in Constitutional AI, described in Section 3.3). For this reason, before presenting the four case studies, we first provide a overview of RL methods and architectures in Section 3.1.
While the case studies discussed in this study focus on the RL approach, the hybrid space of designing AI agents based on moral preferences can, in theory, include a much wider set of methodologies where humans pre-define learning constraints or signals for agents. For example, LLM fine-tuning can instead rely on Supervised Learning based on a set of labels, or Direct Preference Optimisation based on a set of preferences (?), or a mix of approaches, possibly applied in sequence.
@@ 3.1 Background
RL is a method that directly models decision-making, in contrast to prediction or classification - as in Supervised or Unsupervised Learning. In the RL framework, learning occurs incrementally over discrete time steps $t\in T$. At every time step $t$, an agent observes a state of the environment $s^{t}$ and takes an action $a^{t}$ using a policy (a mapping between states and actions). The environment then returns a reward $R^{t+1}$ and a next state $s^{t+1}$. Frequently, RL implementations use the Markov Decision Process framework, in which all necessary information for the agent to make a decision is contained within the last available state.
@@
Over a number of iterations of such interactions with an environment, an agent learns to take actions that maximise the discounted future reward (discounted to make rewards received further in the future less important for the present calculations). Various RL algorithms exist, with two key types being value-based or policy-based. In this paper, we focus on the set of approaches known as model-free, where an agent does not know the model of how the environment may respond to their actions - we believe that this approach applies intuitively to uncertain social situations where other agents whose behaviour is not directly predictable constitute a part of the environment.
@@
A popular value-based RL algorithm is Q-Learning, introduced by ? (?), where an agent learns the ‘value’ of each possible state-action pair $Q(s,a)$. The value represents the estimated discounted cumulative reward an agent may expect to receive by taking this action in this state, and assuming they act optimally from then on. Two versions of Q-learning exist: a tabular version for small state spaces and Deep Q-Network, or DQN (?) for value function approximation in state spaces too large to learn directly. As shown in Figure 3, in tabular Q-learning, the learned value function constitutes a table mapping every state-action pair to the corresponding Q-value. In DQN, on the other hand, the learned representation is a tuned neural network which approximates the state-action value function (?), which takes a state vector representation as input and returns the relative value of taking each possible action as an output. It is this representation that allows DQN agents to generalise knowledge across states, enabling them to choose actions based on past experience even if they have not seen this particular state before.
@@
Figure 3: RL architectures for a hypothetical environment with two possible states and two actions. In tabular Q-learning (1.1) (?), an agent learns a direct mapping from every state to the value $Q$ of taking each possible action in that state. In DQN (1.2) (?), an agent trains a neural network, which takes a representation of the state as input and outputs the associated Q-value for each possible action. In policy-based methods, such as PPO (?), an agent learns a policy model, which maps input states to a probability $P$ of taking each action in that state (in practice, policy-based RL often also uses a separate value model/network for estimating the value $V$ of a given state, not pictured here). The use of neural networks allows for function approximation, which means that not all states need to be observed during training for an agent to learn a (theoretically) general representation.
@@
These representations and the learned value functions can then be used to choose actions according to a certain policy $\pi(s_{t})$ for example, using an $\epsilon$-greedy policy an agent would take a ‘greedy’ action most of the time, i.e., the action with the greatest estimated Q-value available from this state, but act randomly $\epsilon\%$ of the time, sampling uniformly at random from the available action space $A$. The random exploration in this policy allows for new states to be explored by the agent, an essential step in learning via online RL.
In policy-based methods such as PPO (?), an agent learns a policy function $\pi(s_{t})$ (which informs the agent how to act in certain situations to achieve optimal expected reward). In practice, many policy-based methods also require an agent to learn a separate value function $V$ (for estimating the value of certain states for faster convergence). As in value-based methods, these functions can be approximated with neural networks to allow for greater generalisability. Here, the agent learns on-policy, adjusting its strategy as it explores the space of possibilities in the environment.
@@
The problem of moral alignment is fundamental in multi-agent environments, in which agents (and, possibly, humans) with potentially different or competitive goals interact. In these cases, multiple players learn (i.e., iteratively update their value and/or policy estimates) in parallel, and, for each agent, their opponent(s) constitute a part of the environment. The state for a player can then be defined as a history of their opponents’ previous moves. We will now explore four examples of the use of RL for learning morality via explicitly defined principles, illustrating the potential of the hybrid methodology which we argued for in Section 2.3.